OpenStudy (anonymous):

Simplify: 1. square root of 80 2. square root of 72 3. square root of 12

OpenStudy (anonymous):

\[\sqrt{ab} = \sqrt{a}\sqrt{b}\] eg 1. \[\sqrt{80} = \sqrt{16}\sqrt{5} = 4\sqrt{5}\]

OpenStudy (anonymous):

so square root of 72 = square root 12 * square root 6, so IM LOST

OpenStudy (anonymous):

try to get square numbers out of it :) dont worry nearly everyone struggles with it at first

OpenStudy (anonymous):

72 = 4 * 18

OpenStudy (anonymous):

thats a good one because 4 is a square number

OpenStudy (anonymous):

so 2 square root 18?

OpenStudy (anonymous):

yep, can you take out another square number from 18?

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

name some square numbers :D

OpenStudy (anonymous):

see if any go into 18

OpenStudy (anonymous):

My only answers i have is: 6SR2 3SR2 2SR3 0 6 i dont have 4SR18

OpenStudy (anonymous):

2SR18, my bad.

OpenStudy (anonymous):

18 is 2 * 9 isnt it?

OpenStudy (anonymous):

and 9 is a square number

OpenStudy (anonymous):

yeah. so 3*3 =9 so 2SR3?

OpenStudy (anonymous):

well before we had 2SR18 2sqrt(18) = 2sqrt(9)sqrt(2)

OpenStudy (anonymous):

= 2*3*sqrt(2)

OpenStudy (anonymous):

= 6sqrt(2)

OpenStudy (anonymous):

you see?

OpenStudy (anonymous):

Oh. Ok. That makes since. (:

OpenStudy (anonymous):

So with SR of 12...

OpenStudy (anonymous):

have a go with sqrt(12) , its a bit easier

OpenStudy (anonymous):

2*6 3*4

OpenStudy (anonymous):

which one of those has a square number?

OpenStudy (anonymous):

3*4?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

SR 4= 2

OpenStudy (anonymous):

ok, so sqrt(12) = sqrt(4)sqrt(3) = ?

OpenStudy (anonymous):

2SR3?

OpenStudy (anonymous):

yeah! i think you've got it if you're still unsure how about trying to simplify sqrt(20)

OpenStudy (anonymous):

My only answer choices are: 3 6 2SR29 SR6 4SR2

OpenStudy (anonymous):

what for 12?

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

its none of those...

OpenStudy (anonymous):

I have to find the distance between the coordinates: (7,3) and (7,-3) & i got SR12

OpenStudy (anonymous):

oh, nevermind. i added wrong. Its SR of 36 for the answer. & SR of 36 = 6 :)

OpenStudy (anonymous):

ah i see why you want these now :)

OpenStudy (anonymous):

huh? lol

OpenStudy (anonymous):

i thought this was just the question, i didnt realise you were finding distances

OpenStudy (anonymous):

Oh, nuhuh. lol

OpenStudy (anonymous):

Can you help me with something else?

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

im here to help :)

OpenStudy (anonymous):

Determine the length of the major and minor axes of the following ellipses: 1. x^2/16+y^2/9=1

OpenStudy (anonymous):

i havent done ellipses in aaages lol gimme a sec

OpenStudy (anonymous):

lol. ok.

OpenStudy (anonymous):

ok, so this ellipse has centre (0,0) doesnt it?

OpenStudy (anonymous):

Im not sure, I just started learning it.

OpenStudy (anonymous):

ok, i could do a quick run through on ellipses, that way it'll refresh my memory xD

OpenStudy (anonymous):

Thank you so much.

OpenStudy (anonymous):

ok, since you are doing ellipses you have probably seen circle equations before? , of the form \[x^2 + y^2 = r^2\]

OpenStudy (anonymous):

where r is the radius

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

lets think about the circle with equation \[x^2 + y^2 = 1\] it looks like this: (radius is 1) now we're going to stretch it by scale factor 4 in the x - direction so in the equation we replace x by (x/4) \[(\frac{x}{4})^2 + y^2 = 1\]or\[\frac{x^2}{16} + y^2 = 1\] the new graph looks like this: